OpenStudy (anonymous):

There are 10 marbles in a bag, and the marbles are either red or blue. Eric will randomly choose two marbles from the bag, without replacing the first one. If the probability of both marbles' being red is 2 15 , how many BLUE marbles are in the bag?

OpenStudy (anonymous):

I'm guessing you want to say that the probability is 2/15 If that is true, that means that x/10 * x-1/9 = 2/15, or 12 = x*(x-1_ that means that there are 4 red marbles, or 6 blue marbles